Can someone help me on this problem?
I need to find the best buy on pasta sauce (1 qt. is 32 oz).
is it (A) 1 qt. for $$1.29
(B) 45 oz. for $1.59
(C) 48 oz. for $$1.79
(D) 1 qt. 16 oz. for $1.99
Thanks, Lyn

Answer by adamchapman(301)

Work out the price per ounce by dividing the price by the weight.
(A) 1.29/32=0.04 $/oz.
(B)1.59/45=0.035$/oz.
(C)1.79/48=0.037$/oz.
(D)1.99/(32+16)=1.99/48=0.041$/oz.
So The best buy is (B)
I hope this helps
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk